Determine the pressure exerted on a diver at 30 m below the free surface of the sea. Assume a barometric pressure of 101 kPa and the specific gravity of sea water is 1.03. in 2 decimal places


1
Expert's answer
2021-08-12T11:02:22-0400

Atmospheric pressure Pa = 101 kPa

Distance h from the free surface = 30 m

specific gravity of sea water = 1.03

Density of sea water ρ= 1.03×1000=1030 m3/kg

(Specific gravity× density of the reference fluid (water =1000m3/kg) gives density of the sea water)

Pressure exerted on the diver at 30 m below the free surface=

"P_a+ρgh=101000+(1030 \times 9.81 \times 30)=404129 \;Pa"

Therefore, Pressure acting on the diver 30 m below free surface is 404129 Pa or 404.129 kPa